OpenStudy (anonymous):

Find the point at which the line joining the two points (-3, -5) and (7,5) cuts the y-axis.

OpenStudy (mertsj):

First find the equation of the line: 1. Find the slope 2. Choose one of the points and plug into point-slope form 3. Put the equation into slope-intercept form 4. Identify the y-intercept which is the point where the line cuts the y-axis.

OpenStudy (anonymous):

Sorry, I'm not quite sure how to do that... Could you please teach me?

OpenStudy (mertsj):

The definition of slope is: \[m = \frac{y _{2}-y _{1}}{x _{2}-x _{1}}\]

OpenStudy (anonymous):

So, m=5−-5 -5−-3 ?

OpenStudy (mertsj):

\[m=\frac{5-(-5)}{7-(-3)}=\frac{10}{10}=1\]

OpenStudy (mertsj):

Now choose one of the points and plug into: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

So, 10 - 7 = m(10 - -3) ?

OpenStudy (mertsj):

Choose ONE of the points and replace x_1 and y_1 with those coordinates. Do not replace x and y

OpenStudy (mertsj):

Also we now know that m = 1 so replace m with 1

OpenStudy (anonymous):

Okay, so 5−7 = 1(-5−-3) ?

OpenStudy (mertsj):